Tue. 6/21/05
Frequently in the course of my job as an electrical estimator, I must determine the height of existing parking lot light poles from the ground. Given the fact that I cannot physically measure a standing pole from the ground (it could be 35 - 40' tall), shouldn't I be able to accurately determine the height of a light pole from the ground, using trigonometry? For example, can I measure off 20' along the ground and assume a Right Triangle with the pole being the adjacent side to solve the equation?
I am decidedly math handicapped and am looking for an easy and repeatable way to measure light poles from the ground. I am anxiously looking forward to your reply.
Best Regards,
Keith

Found 2 solutions by longjonsilver, cleomenius:
Answer by longjonsilver(2297)

in a right angled triangle, tanX=opp/adj.

If the angle X is 45 degrees, then tan45=1, ie the opposite and the adjacent side lengths are the same.

So, if you have a stick that is the same length as your arm from shoulder to hand and you hold your arm horizontally with the stick at arms length, vertically upwards, then the angle you make looking at the top of the stick is 45 degrees. This is a rough estimate remember.

Now, all you have to do it walk back and forth looking at the light, until the light and the stick fully coincide. Now the triangle you have formed with the light has an angle of 45 degrees too, so the height of the light is the same as the distance it is from your standpoint (plus or minus a bit, since you should get as close to the ground to do it) --> so sit down on the ground and do it that way and you should be fairly close in your estimate. All with no calculator!

Just get a stick the length of your arm. Or better still, start with something you know the height of and stand that distance away and then get a stick and mark off where on the stick, th etop of the height comes, then saw your stick to that height. Then you should have a calibrated stick. WOW! :-)

Jon.

Answer by cleomenius(959)

You should also be able to do this using ratios.
Standing next to the pole, measure your shadow from the sun.
If you know your height to be 72 inches and you shadow is 60 inches,
measure the shadow of the pole.
For example, if the pole shadow is 240 inches, you can set up a ratio
your height = 72
Your shadow = 60
Pole height = x
Pole shadow = 240
Your ratio will be 72/60 = x/240
You can solve for x, and you height will come out to be 288 inches, which will
come out to be 24 feet.
Cleomenius